atoms of CO2 gas decay from upper to lower level with a spread in frequency is about 10kHz. calculate the line shape if the broadening mechanism is found Doppler broadening
1
Expert's answer
2020-06-01T14:18:55-0400

The Gaussian line shape will be


"G=e^{-\text{ln}2\cdot x^2},\\
x=\frac{p-p_0}{w/2},"

where "p, p_0" - position on frequency axis, Hz, "w" - width of the frequency distribution curve where the intensity is half the maximum intensity, Hz.